It takes a girl 36 minutes to deliver the newspapers on her route; however, if her brother helps, it takes them only 20 minutes. How long would it take her brother to deliver the newspapers by himself?
Answer by mananth(16946)

girl takes 36 minutes
1/36 of the job she does in 1 minute
..
together they do do in 20 minutes
1/20 of the job they do in 1 minute
..
1/20 -1/36 = (36-20) / 20*36
=16/20*36
=1/45 of the job he does alone
so he will take 45 minutes to do the job alone.